OpenStudy (anonymous):

Use the properties of logarithms to write the expression as a sum, difference, and/or constant multiple of logarithms. (Assume all variables are positive.) Use the properties of logarithms to write the expression as a sum, difference, and/or constant multiple of logarithms. (Assume all variables are positive.) @Mathematics

OpenStudy (anonymous):

\[\ln \sqrt[3]{t}\]

OpenStudy (anonymous):

\[\ln(xy/z)\]

OpenStudy (anonymous):

= ln t^(1/3) = (1/3) ln t

OpenStudy (anonymous):

ln(xy/z) = ln x + ln y - ln z

OpenStudy (anonymous):

\[\ln((x^{2}-1)/(x ^{3})) \] x>1